OpenStudy (anonymous):

what is the simplification of the expression Square root of 50/9? (drawn below)

OpenStudy (anonymous):



OpenStudy (lgbasallote):

50 = 25 x 2 so turn it into \[\LARGE \sqrt{\frac{25 \times 2}{9}}\] i believe 25 and 9 are perfect squares yes?

OpenStudy (anonymous):

yes they are

OpenStudy (lgbasallote):

so what is the square root of 25?

OpenStudy (anonymous):

5

OpenStudy (lgbasallote):

and the square root of 9?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

so..10 over 3?

OpenStudy (lgbasallote):

therefore it will become \[\frac{5}{3} \sqrt 2\] got it? you cannot multiply 2 because it is not a perfect square..it's still inside the radical

OpenStudy (anonymous):

Oh ok. Thanks!